# The SAGA Survey. III. A Census of 101 Satellite Systems around Milky Way-mass Galaxies

### v2

@@ 5.8 Co-rotating Pairs of Satellites
The upper left panel in Figure 17 shows the fraction of corotating pairs in bins of opening angle. We repeat the analysis with all confirmed satellites passing the velocity cut (blue circles) and with only satellites within 150 kpc in projection to the host galaxies (same velocity cut applies). We calculate the expected errors by repeating our analysis 5,000 times, but each time, we assign each satellite a random position angle sampled uniformly between 0∘ and 360∘. This procedure preserves the satellite radial and velocity distributions but would erase any corotating signal. The 1 interval from this random analysis is shown as the gray band. The two shades of gray band corresponds to the random analyses with two different projected distance cuts. The numbers of pairs in each opening angle are shown in the lower panel.
Overall, the signals are consistent with no corotation (the horizontal line at 0.5). In a few bins, the signal shows 1–2 deviation, which is expected given the number of bins used in the analysis. We do not observe an excess of opposed pairs (small opening angles) that have the same sign of velocity offsets, as recently reported in SDSS data (Gu et al., 2024). The only notable deviation is the rightmost bin: when the satellites in a pair have very similar position angles, there is an excess in the corotation fraction. The excess is more significant (3) in the analysis when we only include satellites within 150 kpc, which hints that the excess comes from pairs with small separation. In Appendix A, a follow-up analysis demonstrates there is a small excess of close satellite pairs, and the satellites in close pairs tend to have a small velocity difference (see Figure A.6).
@@ 6.1 Milky Way in the Cosmological Context
At the high-mass end, the apparent difference in satellite quenched fraction may be explained by the recent accretion of the LMC and the SMC. At the low-mass end, the lack of low-mass star-forming satellites in the Milky Way may be due to earlier satellite infall or the influence of the Local Group (see more discussion in Paper IV). However, the discrepancy in quenched fraction does not have a strong statistical significance considering that the MW is a single host. If we draw a random sample from SAGA’s average quenched fraction using the five most massive MW satellites’ stellar masses, about 16% of the draws will result in the two most massive satellites being star forming and the rest being quenched.
@@ 6.3 Follow-up Studies and Observations
In Section 5.3, we discussed the satellite quenched fraction as a function of stellar mass. We refer readers to Paper IV for a more detailed study on the satellite quenched fraction and star formation rates. Specifically, in Paper IV, we inspect the quenched fraction as a function of projected distance and also compare the satellite quenched fraction between the SAGA and ELVES Surveys. We also study satellite gas-phase metallicities and gas content in Paper IV, repeating our host correlation analysis for the quenched fraction of satellites.
On the theory side, in Section 5.6 we only briefly compare the satellite abundance with the UM-SAGA model. In Paper V, we compare the SAGA DR3 results with the model predictions in detail; generally we find good agreements between SAGA results and simulations. The use of an empirical model such as the UM-SAGA model allows us to study the model ingredients required to reproduce not only the satellite SMFs but also the quenched fraction. We refer readers to Paper V for those discussions.
@@
In addition, we have several planned follow-up observations. Currently, we do not have Hi measurement for all the SAGA satellites (Section 2.5 of Paper IV). Efforts have been made (e.g., Jones et al., 2024) to carry out Hi observation for some SAGA satellites, and we plan to continue this effort for the remaining satellites. Hi measurements for the sample will provide a complete picture on how host–satellite interaction affects the gas content and star formation of the satellites.
@@ 7.2 SAGA DR3 Main Findings
This work was supported by NSF collaborative grants AST-1517148 and AST-1517422 awarded to M.G. and R.H.W. and by Heising-Simons Foundation grant 2019-1402. Additional support was provided by the Kavli Institute for Particle Astrophysics and Cosmology at Stanford and SLAC. Support for Y.-Y.M. during 2019–2022 was in part provided by NASA through the NASA Hubble Fellowship grant no. HST-HF2-51441.001 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555.
@@ Appendix A Supplemental Analyses
In this section we present several supplemental analyses that some readers may find interesting or useful.
@@
Panel A: All galaxies in the photometric catalog.
@@
Panel B: Satellite galaxies (see Section 4.3) with confirmed redshifts. In addition to the density map, individual satellite galaxies are also shown as red points.
@@
Panel C: All galaxies but weighted by (see Section B.1); this is equivalent to probability of finding a SAGA satellite in the color–magnitude plane.
@@
Panel D: Galaxies with literature redshifts. One can observe two dark ridges around and 19.5, which correspond to the survey depths of SDSS and GAMA.
@@
Panel E: Galaxies with redshifts that are first obtained by SAGA. Note that many of the SAGA redshifts are outside the primary targeting region, which is the region below the pink dashed line.
@@
Panel F: This panel shows the fraction of galaxies with confirmed redshifts. In other words, it is the ratio of the combined density shown in Panels D and E to the density shown in Panel A.
@@
The COSMOS and SDSS lines are taken from Nierenberg et al. (2012, Figure 7, top left panel) and Sales et al. (2013, Figure 3, right panel) respectively, and are identical to the version shown in Paper II. The COSMOS hosts are selected from and . The SDSS hosts selected from , and the lower and upper boundaries of the purple band represent the samples of and , respectively. The Local Volume line, on the other hand, is updated to the ELVES Survey result (Carlsten et al., 2022), using only the 14 hosts that are consistent with SAGA host selection (see Section 5.2). The ELVES survey areas and survey incompleteness are taken into account when calculating the average.
@@
For each system property, we split the SAGA systems into four groups, based on the property quartiles. Each group has a quarter of SAGA systems ( 25). Then, we plot the number of satellites as a function of the property (upper row) and the cumulative SMFs for the four groups (lower row). This visualization allows us to understand what might be driving the correction (or lack of). For example, the third column shows that the SAGA hosts that are quenched (lowest sSFR) tend to have more satellites, likely due to those systems’ higher intrinsic masses. On the other hands, the satellite abundance around star-forming hosts do not have strong dependence on the host sSFR, which is consistent with the lack of dependence on the host HI gas mass (fourth column).
To further inspect the correlation between host sSFR and satellite abundance, we show how satellite abundance depends on host sSFR and group mass in Figure A.5. We can observe a weak negative correlation between the host sSFR and system group mass, which contribute to the correlation between host sSFR and satellite abundance. However, at the fixed group mass, the most quenched host galaxies still tend to have more satellites.
To supplement the result presented in Section 5.8, we further examine if close satellite pairs are more likely to have similar velocities. In Figure A.6, we directly investigate how the velocity difference in each pair varies with the projected separation. In bins of projected separation, we show the median and 16th and 84th percentiles of the velocity difference of the two satellites in pair. We also repeat the same random position angle analysis 5,000 times and show the 1 interval (for each statistics) in gray. We can observe that in the leftmost bin (pairs with a projected separation less than 40 kpc), the velocity difference distribution skew toward the lower end. In particular, the median and 16th and 84th percentiles all are lower compared to the random distribution. The lower panel of Figure A.6 also show that there are more close pairs when compared to the random sample (again, at about level). While these results are not statistical significant enough to draw a strong conclusion, they do indicate that the close pairs in SAGA systems contribute to the signal in the last bin of Figure 17.
@@ Appendix B Survey Completeness
the satellite is not detected as a source in the photometric catalog;
@@
the satellite is detected in the photometric catalog but has an underestimated luminosity and fell out from our sample;
@@
We believe the first category is extremely rare and negligible for the purpose of this study. The DESI Legacy Imaging Surveys have a typical depth of ( PSF depth; Dey et al. 2019). Given the SAGA Survey’s magnitude limit of , we expect SAGA’s galaxy sources to be well characterized in the DESI Imaging photometric catalog. As discussed in Section 2.3, we compare our photometric catalog with the SMUDGes catalog (Zaritsky et al., 2023) and found no galaxies missing in our catalog. In addition, we conducted a careful visual inspection of the full SAGA footprint to search for any low surface brightness galaxies that are not identified as sources in the photometric catalog, and we did not find any.
@@
Unlike the first two categories, the third category (sources that we did not obtain confirmed redshifts) can have a significant impact on our science results. Hence, we carefully quantify how the third category contributes to our survey incompleteness. For objects in the third category, we can use the available photometric information to estimate how many satellites we might have missed in the photometric catalog. This estimation is then used as our incompleteness correction as the first two categories have negligible or unimportant contribution to our survey incompleteness. We describe this estimation in detail below.
@@ B.1 Incompleteness Correction Model
First, the photometric properties we choose in this work are , where is the magnitude with a fiber of diameter 1.5′′ in 1′′ Gaussian seeing. In addition, these three photometric properties only enter the model in the specific combination of . The choice of this particular combination will become apparent as we introduce the second modification below. To summarize, the model can be expressed as follows.
Here is the rate (or probability) of finding a SAGA satellite at a given location of the photometric space , and is the maximum value can reach.
@@
In addition, we also only use galaxies brighter than to fit the model. Since the satellite rate is a strong function of the magnitude (), we first find a linear combination of the model parameters such that the satellite rate has the least evolution with respect to magnitude. This procedure results in the specific combination we used here. In other words, we assume that when the satellite rate is modeled to depend on , there will no additional dependence on . We verified this assumption by fitting the model to several magnitude bins.
Once the model is specified, the fitting procedure to obtain is the same as Paper II: we use linear regression to maximize the likelihood function. It is worth noting that we only use the redshift data we collected and compiled to fit the model. There are no extra external constraints such as the form of the SMF. The best-fit model parameters are Panel C of Figure A.1 shows the prediction of the fitted model in the color–magnitude panel.
@@
To calculate the incompleteness correction for a given stellar mass bin, we can simply sum up the incompleteness correction weights of all galaxies in that stellar mass bin. Note that, since we do not have the redshift measurements for these galaxies, the stellar masses used here are calculated assuming these galaxies are at the respective hosts’ distances.
@@ Appendix C Data Tables
In this section we provide the tables for the SAGA DR3 data products. For large tables, only the table schema is shown here; their contents are available in machine-readable and Virtual Observatory (VO) forms.
@@
2 | HOSTID | Unique host identifier, can be used to join satellite table | |
@@
6 | km/s | HRV | Heliocentric recession velocity of host |
@@
9 | mag | KsMag | Absolute Ks-band luminosity, extinction corrected and K-corrected |
@@
13 | mag | gr | g-r color, extinction corrected, derived from DESI Legacy Imaging |
@@
23 | Mpc | sep-massive | Physical distance to nearest massive host |
@@
26 | nsat-G | Number of satellites in Gold sample, with incompleteness correction | |
27 | nsat-GS | Number of satellites in GoldSilver sample, with incompleteness correction | |
28 | nsat-Gc | Number of confirmed satellites in Gold sample | |
29 | nsat-GSc | Number of confirmed satellites in GoldSilver sample | |
@@
Note. — The full contents of this table (101 rows) will be made available in a machine-readable format at publication.
Note. — The full contents of this table (75,704 rows) will be made available in a machine-readable format at publication.
Note. — The full contents of this table (378 rows) will be made available in a machine-readable format at publication.
Note. — The full contents of this table (156 rows) will be made available in a machine-readable format at publication.

### v3

@@ 5.8 Co-rotating Pairs of Satellites
The upper left panel in Figure 17 shows the fraction of corotating pairs in bins of opening angle. We repeat the analysis with all confirmed satellites passing the velocity cut (blue-filled circles) and with only satellites within 150 kpc in projection to the host galaxies (same velocity cut applies). We calculate the expected errors by repeating our analysis 5,000 times, but each time, we assign each satellite a random position angle sampled uniformly between 0∘ and 360∘. This procedure preserves the satellite radial and velocity distributions but would erase any corotating signal. The 1 interval from this random analysis is shown as the gray band. The two shades of gray band corresponds to the random analyses with two different projected distance cuts. The numbers of pairs in each opening angle are shown in the lower panel.
Overall, the signals are consistent with no corotation (the horizontal line at 0.5). In a few bins, the signal shows a 1–2 deviation, which is expected given the number of bins used in the analysis. We do not observe an excess of opposed pairs (small opening angles) that have the same sign of velocity offsets, as recently reported in SDSS data (Gu et al., 2024). The only notable deviation is the rightmost bin: when the satellites in a pair have very similar position angles, there is an excess in the corotation fraction. The excess is more significant (3) in the analysis when we only include satellites within 150 kpc, which hints that the excess comes from pairs with small separation. In Appendix A, a follow-up analysis demonstrates there is a small excess of close satellite pairs, and the satellites in close pairs tend to have a small velocity difference (see Figure A.6).
@@ 6.1 Milky Way in the Cosmological Context
At the high-mass end, the apparent difference in satellite quenched fraction may be explained by the recent accretion of the LMC and the SMC. At the low-mass end, the lack of low-mass star-forming satellites in the Milky Way may be due to earlier satellite infall or the influence of the Local Group (see further discussion in Paper IV). However, the discrepancy in quenched fraction does not have a strong statistical significance, considering that the MW is a single host. If we draw a random sample from SAGA’s average quenched fraction using the five most massive MW satellites’ stellar masses, about 16% of the draws will result in the two most massive satellites being star forming and the rest being quenched.
@@ 6.3 Follow-up Studies and Observations
In Section 5.3, we discussed the satellite quenched fraction as a function of stellar mass. We refer readers to Paper IV for a more detailed study of the satellite quenched fraction and star formation rates. Specifically, in Paper IV, we inspect the quenched fraction as a function of projected distance and also compare the satellite quenched fraction between the SAGA and ELVES surveys. We also study satellite gas-phase metallicities and gas content in Paper IV, repeating our host correlation analysis for the quenched fraction of satellites.
On the theory side, in Section 5.6 we only briefly compare the satellite abundance with the UM-SAGA model. In Paper V, we compare the SAGA DR3 results with the model predictions in detail; generally, we find good agreements between SAGA results and simulations. The use of an empirical model such as the UM-SAGA model allows us to study the model ingredients required to reproduce not only the satellite SMFs but also the quenched fraction. We refer readers to Paper V for those discussions.
@@
In addition, we have several planned follow-up observations. Currently, we do not have Hi measurements for all the SAGA satellites (Section 2.5 of Paper IV). Efforts have been made (e.g., Jones et al., 2024) to carry out Hi observation for some SAGA satellites, and we plan to continue this effort for the remaining satellites. Hi measurements of the sample will provide a complete picture of how host–satellite interaction affects the gas content and star formation of the satellites.
@@ 7.2 SAGA DR3 Main Findings
This work was supported by National Science Foundation (NSF) collaborative grants AST-1517148 and AST-1517422 awarded to M.G. and R.H.W. and by the Heising–Simons Foundation grant 2019-1402. Additional support was provided by the Kavli Institute for Particle Astrophysics and Cosmology at Stanford and SLAC. Support for Y.-Y.M. during 2019–2022 was in part provided by NASA through the NASA Hubble Fellowship grant no. HST-HF2-51441.001 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555.
@@ Appendix A Supplemental Analyses
In this section, we present several supplemental analyses that some readers may find interesting or useful.
@@
Panel (A): All galaxies in the photometric catalog.
@@
Panel (B): Satellite galaxies (see Section 4.3) with confirmed redshifts. In addition to the density map, individual satellite galaxies are also shown as red points.
@@
Panel (C): All galaxies but weighted by (see Section B.1); this is equivalent to probability of finding a SAGA satellite in the color–magnitude plane.
@@
Panel (D): Galaxies with literature redshifts. One can observe two dark ridges around and 19.5, which correspond to the survey depths of SDSS and GAMA.
@@
Panel (E): Galaxies with redshifts that are first obtained by SAGA. Note that many of the SAGA redshifts are outside the primary targeting region, which is the region below the pink dashed line.
@@
Panel (F): This panel shows the fraction of galaxies with confirmed redshifts. In other words, it is the ratio of the combined density shown in Panels (D) and (E) to the density shown in Panel (A).
@@
The COSMOS and SDSS lines are taken from Nierenberg et al. (2012, Figure 7, top left panel) and Sales et al. (2013, Figure 3, right panel), respectively, and are identical to the version shown in Paper II. The COSMOS hosts are selected from and . The SDSS hosts selected from , and the lower and upper boundaries of the purple band represent the samples of and , respectively. The Local Volume line, on the other hand, is updated to the ELVES Survey result (Carlsten et al., 2022), using only the 14 hosts that are consistent with SAGA host selection (see Section 5.2). The ELVES survey areas and survey incompleteness are taken into account when calculating the average.
@@
For each system property, we split the SAGA systems into four groups based on the property quartiles. Each group has a quarter of SAGA systems ( 25). Then, we plot the number of satellites as a function of the property (upper row) and the cumulative SMFs for the four groups (lower row). This visualization allows us to understand what might be driving the correction (or lack of). For example, the third column shows that the SAGA hosts that are quenched (lowest sSFR) tend to have more satellites, likely due to those systems’ higher intrinsic masses. On the other hand, the satellite abundance around star-forming hosts does not have a strong dependence on the host sSFR, which is consistent with the lack of dependence on the host HI gas mass (fourth column).
To further inspect the correlation between host sSFR and satellite abundance, we show how satellite abundance depends on host sSFR and group mass in Figure A.5. We can observe a weak negative correlation between the host sSFR and system group mass, which contributes to the correlation between host sSFR and satellite abundance. However, at the fixed group mass, the most quenched host galaxies still tend to have more satellites.
To supplement the result presented in Section 5.8, we further examine if close satellite pairs are more likely to have similar velocities. In Figure A.6, we directly investigate how the velocity difference in each pair varies with the projected separation. In bins of projected separation, we show the median and 16th and 84th percentiles of the velocity difference of the two satellites in pairs. We also repeat the same random position angle analysis 5,000 times and show the 1 interval (for each statistic) in gray. We can observe that in the leftmost bin (pairs with a projected separation less than 40 kpc), the velocity difference distribution skews toward the lower end. In particular, the median and 16th and 84th percentiles are all lower compared to the random distribution. The lower panel of Figure A.6 also shows that there are more close pairs when compared to the random sample (again, at about level). While these results are not statistically significant enough to draw a strong conclusion, they do indicate that the close pairs in SAGA systems contribute to the signal in the last bin of Figure 17.
@@ Appendix B Survey Completeness
The satellite is not detected as a source in the photometric catalog.
@@
The satellite is detected in the photometric catalog but has an underestimated luminosity and fell out from our sample.
@@
We believe the first category is extremely rare and negligible for the purpose of this study. The DESI Legacy Imaging Surveys have a typical depth of ( point-spread function depth; Dey et al. 2019). Given the SAGA Survey’s magnitude limit of , we expect SAGA’s galaxy sources to be well characterized in the DESI Imaging photometric catalog. As discussed in Section 2.3, we compared our photometric catalog with the SMUDGes catalog (Zaritsky et al., 2023) and found no galaxies missing in our catalog. In addition, we conducted a careful visual inspection of the full SAGA footprint to search for any low surface brightness galaxies that are not identified as sources in the photometric catalog, and we did not find any.
@@
Unlike the first two categories, the third category (sources that we did not obtain confirmed redshifts) can have a significant impact on our science results. Hence, we carefully quantify how the third category contributes to our survey’s incompleteness. For objects in the third category, we can use the available photometric information to estimate how many satellites we might have missed in the photometric catalog. This estimation is then used as our incompleteness correction as the first two categories have negligible or unimportant contribution to our survey incompleteness. We describe this estimation in detail below.
@@ B.1 Incompleteness Correction Model
First, the photometric properties we choose in this work are , where is the magnitude with a fiber of diameter in 1′′ Gaussian seeing. In addition, these three photometric properties only enter the model in the specific combination of . The choice of this particular combination will become apparent as we introduce the second modification below. To summarize, the model can be expressed as follows:
Here, is the rate (or probability) of finding a SAGA satellite at a given location of the photometric space , and is the maximum value can reach.
@@
In addition, we also only use galaxies brighter than to fit the model. Since the satellite rate is a strong function of the magnitude (), we first find a linear combination of the model parameters such that the satellite rate has the least evolution with respect to magnitude. This procedure results in the specific combination we used here. In other words, we assume that when the satellite rate is modeled to depend on , there will be no additional dependence on . We verified this assumption by fitting the model to several magnitude bins.
Once the model is specified, the fitting procedure to obtain is the same as Paper II: we use linear regression to maximize the likelihood function. It is worth noting that we only use the redshift data we collected and compiled to fit the model. There are no extra external constraints such as the form of the SMF. The best-fit model parameters are Panel (C) of Figure A.1 shows the prediction of the fitted model in the color–magnitude panel.
@@
To calculate the incompleteness correction for a given stellar mass bin, we can simply sum up the incompleteness correction weights of all galaxies in that stellar mass bin. Note that since we do not have the redshift measurements for these galaxies, the stellar masses used here are calculated assuming these galaxies are at the respective hosts’ distances.
@@ Appendix C Data Tables
In this section, we provide the tables for the SAGA DR3 data products. For large tables, only the table schema is shown here; their contents are available in a machine-readable format at sagasurvey.org/data and doi.org/10.3847/1538-4357/ad64c4.
@@
2 | HOSTID | Unique host identifier that can be used to join the satellite table | |
@@
6 | km/s | HRV | Heliocentric recession velocity of the host |
@@
9 | mag | KsMag | Absolute -band luminosity, extinction corrected and K-corrected |
@@
13 | mag | gr | color, extinction corrected, derived from DESI Legacy Imaging |
@@
23 | Mpc | sep-massive | Physical distance to the nearest massive host |
@@
26 | nsat-G | Number of satellites in the Gold sample, with incompleteness correction | |
27 | nsat-GS | Number of satellites in the GoldSilver sample, with incompleteness correction | |
28 | nsat-Gc | Number of confirmed satellites in the Gold sample | |
29 | nsat-GSc | Number of confirmed satellites in the GoldSilver sample | |
@@
Note. — The full contents of this table (101 rows) are available in a machine-readable format at sagasurvey.org/data and doi.org/10.3847/1538-4357/ad64c4.
Note. — The full contents of this table (75,704 rows) are available in a machine-readable format at sagasurvey.org/data and doi.org/10.3847/1538-4357/ad64c4.
Note. — The full contents of this table (378 rows) are available in a machine-readable format at sagasurvey.org/data and doi.org/10.3847/1538-4357/ad64c4.
Note. — The full contents of this table (156 rows) are available in a machine-readable format at sagasurvey.org/data and doi.org/10.3847/1538-4357/ad64c4.